# Analysis and experiments of the dissipative Twistcar: direction reversal and asymptotic approximations

### v1

@@ Abstract
Underactuated wheeled vehicles are commonly studied as nonholonomic systems with periodic actuation. Twistcar is a classical example inspired by a riding toy, which has been analyzed using a planar model of a dynamical system with nonholonomic constraints. Most of the previous analyses did not account for energy dissipation due to friction. In this work, we study a theoretical two-link model of the Twistcar while incorporating dissipation due to rolling resistance. We obtain asymptotic expressions for the system’s small-amplitude steady-state periodic dynamics, which reveals the possibility of reversing the direction of motion upon varying the geometric and mass properties of the vehicle. Next, we design and construct a robotic prototype of the Twistcar whose center-of-mass position can be shifted by adding and removing a massive block, enabling demonstration of the Twistcar’s direction reversal phenomenon. We also conduct parameter fitting for the frictional resistance in order to improve agreement with experiments.
@@ I Introduction
Underactuated locomotive systems are systems whose number of degrees-of-freedom (DoF) is greater than the number of its controlled inputs. Such systems often incorporate nonholonomic constraints [1, 2], which restrict permissible velocity directions without decreasing the effective number of DoF of the system. Examples of such systems can be found in terrestrial motion, such as the Snakeboard [3] and the three-wheeled snake [4], or for swimming in fluid, including Purcell’s swimmer [5, 6] and inertia-dominated swimmers [7, 8, 9].
Based on the number of constraints and DoF, those systems can be classified as purely kinematic [10] or mixed kinematic-dynamic systems [11]. Purely kinematic systems, such as the three-linked kinematic snake [12] and Purcell’s swimmer, contain an equal number of nonholonomic constraints and passive (unactuated) DoF of body motion. This property leads to body motion that is governed only by the geometric changes of the robot’s actuated inputs of shape variables. When the number of constraints is smaller, changes in the system’s momentum influence its movement alongside the kinematic constraints [13]. A classical example of a mixed system is Chaplygin’s sleigh [14], and its extensions with oscillatory actuation, such as the Chaplygin’s beanie [15] and Chaplygin sleigh with a moving mass [16]. Other examples consist of wheeled rolling board toys such as the Snakeboard and Roller-racer [17], which is also known in the literature as the Landfish [18], Land-shark [19] or Twistcar [20] - all are renamed variants of the same toy, originally called Roller-racer.
@@
This theoretically-unbounded behavior arises from the absence of energy dissipation in the system, since the forces maintaining the nonholonomic no-skid constraints do not perform mechanical work. Several methods have been considered to incorporate the wheels’ skid, a behavior also shown in the experiments, into the mathematical model: Some works considered replacing the nonholonomic constraints with a Coulomb’s friction model [25, 26, 22]. Although physically plausible, this hybrid model is numerically sensitive and cannot be studied using analytical tools such as perturbation expansion [20], due to its non-smoothness. The work [19], considered a skid-angle model [27, 28] which mimics skid velocities without adding dissipation to the system. While this model results in smooth dynamics, it still leads to unbounded diverging oscillations due to a lack of dissipation. Another approach introduces dissipation to the system while keeping the nonholonomic no-skid constraints. Recent work [24] suggests using viscous dissipation in rolling resistance of the wheels, and another work used a similar approach on a Snakeboard model [29]. This approach leads to bounded behavior with smooth dynamics.
In this work, we revisit the Twistcar dynamics using the latter assumption of viscous dissipation due to wheels’ rolling resistance, while preserving the nonholonomic no-skid constraints at the lateral directions. We conduct a small-amplitude analysis of the model using perturbation expansion in order to obtain approximate explicit expressions which are compared to numerical simulations, with excellent agreement for small actuation amplitudes. In addition, we design and construct a robotic prototype of the Twistcar with an open-loop controller of the steering joint’s angle. The robot has a modular structure which enables changing the robot’s geometry and mass distribution. The perturbation analysis, simulations, and experiments results in a bounded oscillatory motion. Using the asymptotic solution, we obtain explicit conditions for direction reversal due to adding or removing a massive block onto the robot’s body. The theoretical predictions are corroborated by the motion of our modular robot, thus demonstrating direction reversal in actual motion experiments.
@@ II Problem formulation
We now present formulation of the Twistcar’s dynamic equations of motion, a planar model analogous to that described by [22] as depicted in figure 1. The vehicle has two rigid links, each characterized by a total length $l_{i}$, a center of mass (CoM) located at a distance $b_{i}$ from point $P_{i}$, a mass $m_{i}$, and a moment of inertia $J_{i}$ about its CoM. Two wheels are positioned symmetrically at a distance $d$ from point $P_{1}$, while a third wheel is mounted at the apex of link 2. We introduce a point-mass $m_{0}$, located at a distance $b_{0}$ from point $P_{1}$. We choose the following generalized coordinates:
@@
To enforce the no-skid constraints, we express the slip velocities in terms of the generalized coordinates, arrange them in matrix form, and set them to zero. We give the non-holonomic constraints in (3).
@@
$\textbf{W}=\begin{bmatrix}-\sin{\theta}&\cos{\theta}&0&0\\ -\sin{\left(\theta+\phi\right)}&\cos{\left(\theta+\phi\right)}&l_{2}+l_{1}\cos{\phi}&l_{2}\end{bmatrix}$
@@
$\textbf{v}_{\parallel,i}=\mathbb{J}_{i}(\textbf{q})\dot{\textbf{q}}$ (4)
@@
To derive the constrained dynamic equations of motion, we use the Euler–Lagrange method [30], and include the nonholonomic forces acting on the model. We use a planar model with constant potential energy to obtain (7).
@@ III Simulations
In this section, we reformulate the dynamic equations in order to conduct numerical analysis. It allows us to obtain time solutions for different model parameters, showing the bounded behavior of the model and the direction reversal phenomenon.
Figure 2: Numeric simulation results for the dissipative and non-dissipative models, using the parameters in (12), (13). (a) The vehicle’s speed in the world frame. (b) The angle of Link 1 about the average direction of progression. (c) Nonholonomic no-skid constraint forces.
@@
Using (11), we numerically solve the first three equations to find the time solutions of the passive coordinates $\textbf{q}_{p}(t)$. All simulations were conducted using MATLAB’s ode45 function. To obtain the input torque $\tau$ and the constraint forces $\boldsymbol{\Lambda}$, we substitute $\textbf{q}_{p}(t)$ into the last three equations in (11).
To determine the movement direction and velocity of the vehicle, we examine its velocity along the average direction of progression in world frame. We first simulate under zero initial conditions and obtain the mean value of the steering angle in steady-state $\overline{\theta}$. By substituting $\overline{\theta}$ from the orientation angle’s time result, we obtain a solution that oscillates around the world frame’s $\hat{x}$ axis.
Simulations were conducted using (11), with an input angle as in (12).
@@
$m_{0}=0,\quad$ $m_{1}=1[Kg],\quad m_{2}=0.3[Kg],\quad c=0.5[Kg/s],$
@@
In figure 2 we show numerical time solutions for both dissipative and non-dissipative models. For the non-dissipative model, we use the same formulation and parameters as the dissipative, while taking $c=0$. In contrast to the nonphysical unbounded behavior of the non-dissipative model, the dissipative model has a bounded oscillatory steady-state behavior.
@@
In figure 3 we show how changing the mass distribution of the vehicle influences its movement direction. Without the point-mass, the vehicle moves backward, toward link 1. By adding a point-mass $m_{0}$, the vehicle’s speed changes from negative to positive, meaning the vehicle moves forward toward link 2. Using this result, we design an experimental setup that captures the direction reversal phenomenon, as shown below in Section V-C.
@@ IV Asymptotic analysis
Where $v_{\parallel}$ and $v_{\perp}$ are the velocity components of point $P_{1}$ along $\textbf{e}_{1}^{\prime}$ and $\textbf{e}_{2}^{\prime}$ directions, i.e. parallel and perpendicular the the vehicle’s body.
@@
This defines two relations between the body-frame velocities. The first directly dictates that the lateral velocity is zero. We use the second row of (16) in order to eliminate $\dot{\theta}$, and express the equations of motion using a reduced vector of unconstrained velocities, defined as:
@@
Note that from (21) on, we refer to the forward speed of the vehicle $v_{\parallel}$ as $v$. We use (21) to solve for $v$.
@@
We define the characteristic time of the problem as $t_{c}=m_{1}/c$, and denote $\tilde{t}=t/t_{c}$ and $\tilde{\omega}=\omega t_{c}$. The nondimensional time-dependent variables are:
@@ IV-A Symmetric input
We also expand $\tilde{f}$ as a power series in $\varepsilon$. Under zero initial conditions, the leading and first-order solutions are nullified $v_{0}(t)=v_{1}(t)\equiv 0$, and $v(t)$ is an even power series. The second-order ODE is in the form of:
@@
In figure 4(a), we present the nondimensional analytic solution for the forward speed $v(t)$ overlaid with the numeric solution, both obtained with the parameter values as in (13). We check the fit quality between the two solutions in figure 4(c), plotting the mean forward speed in steady-state for different input amplitudes on a log-log scale. The analytic solution is a linear line with a slope of 2 due to its leading order $\varepsilon^{2}$, and we see a good alignment between this line and the numeric solution for $\varepsilon\ll 1$ and even for input amplitudes which are closer to 1.
Next, we examine the conditions for direction reversal. We note that the sign of the mean value of the solution in steady-state, $b_{1}$, dictates the direction of movement of the vehicle. In (29) we show $\xi$, the condition for direction reversal under the assumption $\beta_{2}=1/2$. $\xi$ is positively proportional to $b_{1}$, consisting of the terms that dictate the sign of $a_{1}$. When $\xi>0$, the mean velocity of the vehicle is positive, and when $\xi<0$, the mean velocity of the vehicle is negative.
@@
We compare our results to the non-dissipative model in [22], in which the direction reversal condition is determined by the mean acceleration of the vehicle. Under the same assumption of $\beta_{2}=1/2$, we get the same condition for direction reversal.
@@ IV-B Asymmetric input
We consider an input with small deviation $\phi_{0}\ll 1$ from the symmetric case. We write (21) in a nondimensional form and expand $v(t)$ as in (25). As a result, we get the following solution:
@@
Figure 5: Comparison between numeric simulation and analytic approximation using (13) and $\varepsilon=\pi/10$. (a) Nondimensional forward speed comparison with $\phi_{0}=5^{\circ}$. (b) Dimensional mean curvature of the robot’s path, the dash-dotted line is the value of the input angle’s amplitude.
The correction term in (33) oscillates around zero at steady-state, meaning it does not affect the direction reversal condition. In addition, it oscillates at the input frequency, in contrast to $v_{2}(t)$ which oscillates at twice this frequency. In figure 5(a), we show a comparison between the numeric simulation and analytic approximation of the forward speed $v(t)$ using the parameters in (13), while incorporating an asymmetric mean input angle.
@@
We use (30) to find the orientation angle’s angular velocity, substituting the velocity as in (32) and expanding as a power series in $\varepsilon$. We show the steady-state solution of the angular velocity in (34). This solution contains a mean value that appears due to the asymmetric input. This constant change in the mean orientation angle of the vehicle leads to a motion along a mean circular arc.
@@
In figure 5(b), we present the curvature of the robot’s mean path, depending on the mean input angle. Assuming that the input angle’s mean value is up to the order of its amplitude, we get that the approximated curvature is:
$\rho=\frac{\overline{\theta_{ss}^{\prime}}}{\overline{v}}=\frac{\varepsilon^{2}\phi_{0}d_{1}+\mathcal{O}\left({\phi_{0}}^{2}\varepsilon\right)}{\varepsilon^{2}b_{1}+\mathcal{O}\left({\phi_{0}}^{2}\varepsilon,\phi_{0}\varepsilon^{2},\varepsilon^{3}\right)}=\frac{d_{1}}{b_{1}}\phi_{0}+\ldots$ (35)
The simulation results show that the curvature is linearly proportional to the mean angle value for an input angle’s mean value smaller than its amplitude. For a larger mean value, (35) is no longer satisfied having a nonlinear relation.

### v2

@@ Abstract
Underactuated wheeled vehicles are commonly studied as nonholonomic systems with periodic actuation. Twistcar is a classical example inspired by a riding toy, which has been analyzed using a planar model of a dynamical system with nonholonomic constraints. Most of the previous analyses did not account for energy dissipation due to frictional resistance. In this work, we study a theoretical two-link model of the Twistcar while incorporating dissipation due to rolling resistance. We obtain asymptotic expressions for the system’s small-amplitude steady-state periodic dynamics, which reveals the possibility of reversing the direction of motion upon varying the geometric and mass properties of the vehicle. Next, we design and construct a robotic prototype of the Twistcar whose center-of-mass position can be shifted by adding and removing a massive block, enabling experimental demonstration of the Twistcar’s direction reversal phenomenon. We also conduct parameter fitting for the frictional resistance in order to improve agreement with experiments.
@@ I Introduction
Underactuated locomotive systems are systems whose number of degrees-of-freedom (DoF) is greater than the number of its controlled inputs. Such systems often incorporate nonholonomic constraints [1, 2], which restrict permissible velocity directions without decreasing the effective number of the system’s DoF. Examples of such systems can be found in terrestrial motion, such as the Snakeboard [3] and the three-wheeled snake [4], or for swimming in fluid, including Purcell’s swimmer [5, 6] and inertia-dominated swimmers [7, 8, 9].
Based on the number of constraints and DoF, those systems can be classified as purely kinematic [10] or mixed kinematic-dynamic systems [11]. Purely kinematic systems, such as the three-linked kinematic snake [12] and Purcell’s swimmer, contain an equal number of nonholonomic constraints and passive (unactuated) DoF of body motion. This property leads to body motion that is governed only by the geometric changes of the robot’s actuated inputs of shape variables. When the number of constraints is smaller, changes in the system’s momentum influence its movement alongside the kinematic constraints [13]. A classical example of a mixed system is Chaplygin’s sleigh [14], and its extensions with oscillatory actuation, such as the Chaplygin’s beanie [15] and Chaplygin sleigh with a moving mass [16]. Other examples consist of wheeled rolling board toys such as the Snakeboard and Roller-racer [17, 18], which is also known in the literature as the Landfish [19], Land-shark [20] or Twistcar [21] - all are renamed variants of the same toy, originally called Roller-racer.
@@
This theoretically-unbounded behavior arises from the absence of energy dissipation in the system, since the forces maintaining the nonholonomic no-skid constraints do not perform mechanical work. Several methods have been considered to incorporate the wheels’ skid, a behavior also shown in the experiments, into the mathematical model: Some works considered replacing the nonholonomic constraints with a Coulomb’s friction model [26, 27, 23]. Although physically plausible, this hybrid model is numerically sensitive and cannot be studied using analytical tools such as perturbation expansion [21], due to its non-smoothness. The work [20], considered a skid-angle model [28, 29] which mimics skid velocities without adding dissipation to the system. While this model results in smooth dynamics, it still leads to unbounded diverging oscillations due to a lack of dissipation. Another approach introduces dissipation to the system while keeping the nonholonomic no-skid constraints. Recent work [25, 18] suggests using viscous dissipation in rolling resistance of the wheels, and another work used a similar approach on a Snakeboard model [30]. This approach leads to bounded behavior with smooth dynamics.
In this work, we revisit the Twistcar dynamics using the latter assumption of viscous dissipation due to wheels’ rolling resistance, while preserving the nonholonomic no-skid constraints at the lateral directions. We conduct a small-amplitude analysis of the model using perturbation expansion in order to obtain approximate explicit expressions which are compared to numerical simulations, with excellent agreement for small actuation amplitudes. In addition, we design and construct a robotic prototype of the Twistcar with an open-loop controller of the steering joint’s angle. The robot has a modular structure which enables changing the robot’s geometry and mass distribution. The perturbation analysis, simulations, and experiments, all result in a bounded oscillatory motion. Using the asymptotic solution, we obtain explicit conditions for direction reversal due to adding or removing a massive block onto the robot’s body. The theoretical predictions are corroborated by the motion of our modular robot, thus demonstrating direction reversal in actual motion experiments.
@@ II Problem formulation
We now present formulation of the Twistcar’s dynamic equations of motion, a planar model analogous to that described by [23] as depicted in Figure 1. The vehicle has two rigid links, each characterized by a total length $l_{i}$, a center of mass (CoM) located at a distance $b_{i}$ from point $P_{i}$, a mass $m_{i}$, and a moment of inertia $J_{i}$ about its CoM. Two wheels are positioned symmetrically at a lateral distance $d$ from point $P_{1}$, while a third wheel is mounted at the apex of link 2. We introduce a point-mass $m_{0}$, located at a distance $b_{0}$ from point $P_{1}$. We choose the following generalized coordinates:
@@
To enforce the no-skid constraints, we express the wheels’ skid velocities in terms of the generalized coordinates:
$\textbf{v}_{\perp,i}=\mathbb{J}_{\perp,i}(\textbf{q})\dot{\textbf{q}},\quad i=1,2$ (3)
Using the Jacobian notation, we define the nonholonomic no-slip constraints in (4), by setting the slip velocities to zero:
@@
$\textbf{W}=\begin{bmatrix}-\sin{\theta}&\cos{\theta}&0&0\\ -\sin{\left(\theta+\phi\right)}&\cos{\left(\theta+\phi\right)}&l_{2}+l_{1}\cos{\phi}&l_{2}\end{bmatrix}$ (5)
@@
$\textbf{v}_{\parallel,i}=\mathbb{J}_{i}(\textbf{q})\dot{\textbf{q}},\quad i=1,2,3$ (6)
@@
To derive the constrained dynamic equations of motion, we use the Euler–Lagrange method [31], and include the nonholonomic forces acting on the model. We define the total kinetic energy of the problem $T\left(\textbf{q},\dot{\textbf{q}}\right)$ and use a planar model with constant potential energy to obtain (9).
@@ III Simulations
In this section, we reformulate the dynamic equations in order to conduct numerical analysis. It allows us to obtain the system’s solution in time for different model parameters, showing the bounded behavior of the model and the direction reversal phenomenon.
Figure 2: Numeric simulation results for the dissipative and non-dissipative models, using the parameters in (14), (15). (a) The vehicle’s speed $\dot{x}(t)$ in the world frame. (b) The angle of Link 1 relative to the average direction of progression. (c) Nonholonomic no-skid constraint forces.
@@
Using (13), we numerically solve the first three equations to find the time solutions of the passive coordinates $\textbf{q}_{p}(t)$. All simulations were conducted using MATLAB’s ode45 function. To obtain the input torque $\tau(t)$ and the constraint forces $\boldsymbol{\Lambda}(t)$, we substitute the solution $\textbf{q}_{p}(t)$ into the last three equations in (13).
To determine the movement direction and velocity of the vehicle, we examine its velocity along the average direction of progression in the world frame. We first simulate under zero initial conditions and obtain the mean value of the orientation angle in steady-state $\overline{\theta}$. By subtracting $\overline{\theta}$ from $\theta(t)$, we obtain a solution such that it oscillates around zero, and the vehicle’s net displacement is along the world frame’s $\hat{x}$ axis.
Simulations were conducted using (13), with an input angle as in (2), for:
@@
$m_{0}=0,\quad$ $m_{1}=1[Kg],\quad m_{2}=0.3[Kg],\quad c=0.5[N\cdot s/m],$
@@
In Figure 2 we show numerical time solutions for both dissipative and non-dissipative models. For the non-dissipative model, we use the same formulation and parameters, except for taking $c=0$. In contrast to the unphysical unbounded behavior of the non-dissipative model, the dissipative model has a bounded oscillatory steady-state behavior.
@@
In Figure 3 we show how changing the mass distribution of the vehicle influences its movement direction. Without the point-mass $m_{0}=0$, the vehicle moves backward $v(t)<0$, toward link 1. By adding a point-mass $m_{0}$, the vehicle’s speed changes from negative to positive, meaning the vehicle moves forward toward link 2. Using this result, we design an experimental setup that captures the direction reversal phenomenon, as shown below in Section V-C.
@@ IV Asymptotic analysis
Where $v_{\parallel}$ and $v_{\perp}$ are the velocity components of point $P_{1}$ along $\textbf{e}_{1}^{\prime}$ and $\textbf{e}_{2}^{\prime}$ directions, i.e. parallel and perpendicular to the vehicle’s body.
@@
This defines two relations between the body-frame velocities. The first directly dictates that the lateral velocity $v_{\perp}$ is zero. We use the second row of (18) in order to eliminate $\dot{\theta}$, and express the equations of motion using a reduced vector of unconstrained velocities, defined as:
@@
Note that from (23) on, we refer to the forward speed of the vehicle $v_{\parallel}$ as $v$. We use (23) to solve for $v(t)$ under prescribed input $\phi(t)$.
@@
We define the characteristic time of the problem $t_{c}=m_{1}/c$, and denote $\tilde{t}=t/t_{c}$ and $\tilde{\omega}=\omega t_{c}$. The nondimensional time-dependent variables are:
@@ IV-A Symmetric input
After substituting (28) into (27), We expand $\tilde{f}$ as a power series in $\varepsilon$ and arrange the expression in powers of $\varepsilon$. Under zero initial conditions, the leading and first-order solutions are nullified $v_{0}(t)=v_{1}(t)\equiv 0$, and $v(t)$ is an even-power series. The second-order ODE is in the form of:
@@
In Figure 4(a), we present the nondimensional analytic solution for the forward speed $v(t)$ overlaid with the numeric solution, both obtained with the parameter values as in (14) and (15). We check the fit quality between the two solutions in Figure 4(c), plotting the mean forward speed in steady-state for different input amplitudes on a log-log scale. The analytic solution is a linear line with a slope of 2 due to its leading order $\varepsilon^{2}$, and we see a good alignment between this line and the numeric solution for $\varepsilon\ll 1$ and even for input amplitudes which are closer to 1.
Next, we examine the conditions for direction reversal. We note that the sign of the mean value of the solution $v(t)$ in steady-state, $b_{1}$, dictates the direction of movement of the vehicle. In (31) we show $\xi$, the condition for direction reversal under the simplifying assumption $\beta_{2}=1/2$. $\xi$ is positively proportional to $b_{1}$, consisting of the terms that dictate its sign. When $\xi>0$, the mean velocity of the vehicle is positive, and when $\xi<0$, the mean velocity of the vehicle is negative.
@@
We compare our results to the non-dissipative model in [23], in which the direction reversal condition is determined by the mean acceleration of the vehicle. Under the same assumption of $\beta_{2}=1/2$, we obtain the same condition for direction reversal.
@@ IV-B Asymmetric input
We consider an input with small deviation $\phi_{0}\ll 1$ from the symmetric case. We write (23) in a nondimensional form and expand $v(t)$ as in (27). As a result, we obtain the following solution, assuming $\phi_{0}\ll\varepsilon$:
@@
Figure 5: Comparison between numeric simulation and analytic approximation using (15) and $\varepsilon=\pi/10[rad]$. (a) Nondimensional forward speed comparison with asymmetry $\phi_{0}=5^{\circ}$. (b) Dimensional mean curvature of the robot’s path, the dashed-dotted line is the value of the input angle’s amplitude $\varepsilon$.
The correction term in (35) oscillates around zero at steady-state, meaning it does not affect the direction reversal condition. In addition, it oscillates at the input frequency, in contrast to $v_{2}(t)$ which oscillates at twice this frequency. In Figure 5(a), we show a comparison between the numeric simulation and analytic approximation of the forward speed $v(t)$ using the parameters in (15), while incorporating an asymmetric mean input angle of $\phi_{0}=\pi/36[rad]$.
@@
We use (32) to find the orientation angle’s velocity $\theta^{\prime}(t)$ by substituting the velocity as in (34) and expanding as a power series in $\varepsilon$. We show the steady-state solution of the angular velocity in (36). This solution contains a mean value that appears due to the asymmetric input. This constant change in the mean orientation angle of the vehicle leads to a motion along a mean circular arc.
@@
We derive the approximate curvature of the robot’s mean path using the asymmetric analytic solutions in (34) and (36):
$\rho=\frac{\overline{\theta_{ss}^{\prime}}}{\overline{v}}=\frac{\varepsilon^{2}\phi_{0}d_{1}+\mathcal{O}\left({\phi_{0}}^{2}\varepsilon\right)}{\varepsilon^{2}b_{1}+\mathcal{O}\left({\phi_{0}}^{2}\varepsilon,\phi_{0}\varepsilon^{2},\varepsilon^{3}\right)}\approx\frac{d_{1}}{b_{1}}\phi_{0}$ (37)
In Figure 5(b), we present a comparison between the numeric and approximate curvatures. This comparison highlights that under the assumption $\phi_{0}\ll\varepsilon$, left of the dashed-dotted vertical line, the curvature is linearly proportional to the mean angle value. Otherwise, the discrepancy from (37) becomes significant.